OpenStudy (anonymous):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen?

OpenStudy (debbieg):

You need to find the volume of the paint tray. How would you do that? (keeping in mind, you have inches and cm both involved - mixing apples and orange, lol - so you'll want to convert one).

OpenStudy (debbieg):

Since the volume of the gallon is given in inches, it would probably makes sense to convert the depth of the tray to inches.

OpenStudy (anonymous):

i need help on this question to

OpenStudy (debbieg):

OK, I explained above what needs to be done. You need to compute the volume of the tray. V=LxWxH You have all of those, except 2 are in inches and 1 in cm, so convert that one to inches (since the volume of the gallon of paint is given in cubic inches). Once you know the volume of the tray, you can compare that to the volume of the gallon.

OpenStudy (anonymous):

so it would over fill the tray

OpenStudy (debbieg):

What did you get for the cubic volume of the tray?

OpenStudy (anonymous):

240 rounded

OpenStudy (debbieg):

How did you arrive at that? If the tray holds 240 cubic inches, will the gallon overflow?

OpenStudy (anonymous):

2.54=1 inch. so 2.54*2=5.08 so....

OpenStudy (debbieg):

Ok, that isn't how you do a conversion, so let's talk about that. conversions are easy - don't make them too hard! :) All you do to convert units is to muliply by the correct CONVERSION FACTOR. The conversion factor must have something in the num'r that is = to what's in the den'r, but just different units. What goes in the num'r and what in the den'r, depends on what units you are STARTING WITH and what units you want to END WITH. So if I have, for example, hours, and I want minutes, I need to use the fact that 1 hour = 60 minutes (see, they are equal?). So my conversion factor is EITHER \(\dfrac{ 60~\min }{ 1~hour }\) or \(\dfrac{ 1~hour }{ 60~\min }\) Which I use depend on whether I'm converting from hours to minutes (so I want HOURS to cancel, thus I need HOURS in the den'r (bottom) and minutes in the num'r (top)). Whatever goes in the TOP is what you'll be LEFT WITH. Whatever goes in the BOTTOM is what will CANCEL. Here we have that 1 inch = 2.54 cm, so the conversion factor is either: \(\dfrac{ 2.54~cm }{ 1~inch }\) or \(\dfrac{ 1~inch }{ 2.54~cm }\) You want to go FROM cm TO inches, to it will be the SECOND option, right?

OpenStudy (debbieg):

So the conversion is: \(4~cm\cdot\dfrac{ 1~inch }{ 2.54~cm }=4~\cancel{cm}\cdot\dfrac{ 1~inch }{ 2.54~\cancel{cm} }=\dfrac{ 4 }{ 2.54}~inches\)

OpenStudy (debbieg):

Now just put that into you volume calculation.

OpenStudy (anonymous):

so you divide 2.54 into 4 right

OpenStudy (debbieg):

yes. just enter it on you calculator: 4/2.54

OpenStudy (debbieg):

*your

OpenStudy (anonymous):

on my calculator i got 1.57

OpenStudy (debbieg):

ok, good. That's the depth in inches. Now multiply that by L*W

OpenStudy (debbieg):

You can even do the whole thing as one entry: you want L*W*H (h is depth here) so: \(10~in\cdot 14~in\cdot\dfrac{ 4 }{ 2.54}~in=???~in^3\)

OpenStudy (debbieg):

(but it's also fine to do the conversion first, then take that product with the other two dimensions)

OpenStudy (anonymous):

i restarted my exam so the question changed to this Alaine has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 12 inches long, and 5 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 369 cm3. The paint will not fill the tray by 5.22 in3. The paint will fill the tray exactly. The paint will overfill the tray by 5.22 in3. So actually the dept would be 1.96

OpenStudy (anonymous):

10 * 12 * 1.96 = 235.2

OpenStudy (debbieg):

Well the process is just as discussed above so you should be able to do this one on your own. If this is an exam, I highly doubt that you are supposed to getting assistance from other people.

OpenStudy (anonymous):

all assessments are labeled as exam, im home-school, its not an actual i can redo it you cant redo other exams. This test has live lesson available for it but it isnt tell Wednesday and i need to get it done today to get back on pace but thanks for the help though